OpenStudy (anonymous):

Simplify. 20. (n^2 – 3n + 1) – (4n^2 + n – 5) 21.4y^2z(2y + 3z – 5) 22.(x – 8)(x + 5) Please help, I need this by the end of the day, or may not pass. Please :(

OpenStudy (anonymous):

20. -3 n^2 - 4 n+6

OpenStudy (anonymous):

Thank you so much, still need 21 and 22 though guys!

OpenStudy (anonymous):

21.4y^2z(2y + 3z – 5) in this is it y square only or is it 4y raise to 2z(2y+3z-5) 22.(x – 8)(x + 5) Use this method -> (x-a)(x+b) = x(x+b)-a(x+b) =x^2 + xb - ax -ab Try solving